### v3

# Non-Intrusive Graph-Based Bot Detection for E-Commerce Using Inductive Graph Neural Networks

Abstract
Malicious bots increasingly abuse e-commerce services (scraping, inventory hoarding, fraud) while evading conventional defenses. IP/rule blocking is brittle under proxies and botnets, and CAPTCHAs add friction yet are increasingly bypassed. We propose a non-intrusive, deployment-friendly framework that models session–URL interactions as a bipartite graph and applies an inductive GNN (GraphSAGE) to classify session nodes. By combining topology with lightweight behavioral and URL semantics, the model detects “feature-normal” automation. On real-world traffic with high-confidence bot labels, GraphSAGE outperforms a strong session-feature MLP baseline in AUC and F1. We further test robustness under adversarial edge perturbations and inductive generalization in a cold-start simulation; performance degrades modestly under mild perturbation and remains strong on unseen nodes, supporting real-time deployment without client-side instrumentation.
I Introduction
E-commerce platforms face persistent automated abuse (scraping, inventory hoarding, fraud) from bots that increasingly mimic human browsing. Common defenses are either brittle (IP/user-agent blocking under proxy rotation) or intrusive (CAPTCHAs), motivating passive, high-precision detection from backend telemetry.
Graphs compactly capture relationships across sessions and content, enabling detection beyond per-session aggregates and improving robustness to evolving behavior (e.g., BotChase [1]). We construct a bipartite session–URL interaction graph from standard logs and apply an inductive GNN, GraphSAGE [2], to classify session nodes. Inductive learning supports cold-start scoring for new sessions and new pages; we ask whether this non-intrusive graph signal improves accuracy while remaining robust to graph perturbations and temporal shift.
Our main contributions are summarized as follows:
  • Non-intrusive graph formulation: We cast bot detection as learning on a session–URL graph built from standard server logs, avoiding CAPTCHAs and client-side instrumentation.
  • Inductive GraphSAGE with lightweight features: We use GraphSAGE to aggregate session and URL attributes, enabling immediate scoring of unseen session/URL nodes via learned neighborhood aggregation.
  • Robustness + deployability evaluation: We report gains over a session-feature MLP baseline and study adversarial edge perturbations and cold-start generalization, showing modest degradation under moderate perturbation and strong performance on unseen nodes.
The remainder of this paper reviews related work (Section II), presents the method (Section III), reports experiments including perturbation and cold-start simulations (Section IV), analyzes results and deployment considerations (Section V), and concludes (Section VI).
II Related Work
II-A Bot Detection in Web and E-commerce
Traditional web bot defenses rely on static indicators or challenges, but modern bots evade them via proxies and automation. Recent work thus emphasizes behavior-driven ML detection to improve adaptability while reducing friction [6, 4]; we follow this passive, high-precision direction.
II-B Graph-Based Fraud and Bot Detection
Relational structure is often diagnostic in security and fraud. BotChase [1] shows improved robustness via graph learning, and GNNs are increasingly used in fraud workflows where relational context improves detection and operational utility [3], motivating session–content graphs for e-commerce.
II-C Graph Neural Networks in Anomaly Detection
GNNs are widely used for graph anomaly detection [5], but production graphs are dynamic and require inductive handling of unseen nodes. GraphSAGE [2] learns feature-driven aggregators that generalize beyond the training graph, fitting live e-commerce traffic.
III Method
Data Ingestion (Raw Server Logs) Feature Extraction (Session & URL Stats) Graph Construction (Bipartite Session-URL) GraphSAGE Model (Inductive Learning) *Filtering Static Resources *Score: Bot vs Human
Figure 1: Overall Architecture. The proposed framework transforms raw logs into behavioral features, constructs a filtered bipartite interaction graph, and utilizes an inductive GraphSAGE model for real-time bot detection.
Our method comprises graph construction, feature design, model architecture, and training/inference. Figure 1 summarizes the end-to-end pipeline.
III-A Graph Construction
We construct a heterogeneous bipartite graph $G=(V,E)$ that links sessions to accessed URLs.
  • Session nodes: Each node is a user session (a sequence of requests/actions within a time window).
  • Content/URL nodes: Each node is a unique page/resource (e.g., product, category, search).
An edge $e\in E$ connects a session node to a URL node if the session accessed the URL. For message passing we treat edges as undirected and unweighted; repeated requests are captured by session features rather than edge multiplicity.
[…]
Figure 2: Illustration of the bipartite session–URL interaction graph used in this work. Nodes represent sessions and accessed URLs, and edges indicate page visits.
Figure 2 illustrates the bipartite session–URL graph used throughout this work.
Motivation: Legitimate sessions follow common navigation patterns, while bots often induce atypical connectivity (broad coverage, rare-page combinations, or coordinated targeting). The graph supports “suspicion by association” via shared URL neighborhoods [3].
We parse logs, map requests to sessions, add nodes for unique sessions/URLs, and update edges online.
CSS
(a) Raw Graph (Noisy Hubs)
P1P2
(b) Refined Graph (Clean)
Figure 3: Graph refinement process. (a) Raw graphs are dominated by static resource hubs (e.g., CSS). (b) Filtering yields meaningful clusters.
Figure 3 shows why we filter static-resource hubs to obtain a cleaner graph for message passing.
III-B Feature Design
We assign lightweight feature vectors to session and URL nodes.
Figure 4: Distribution of representative session-level behavioral features for human and bot sessions. Although individual features exhibit partial overlap, the distributions reveal systematic differences, motivating the use of relational graph modeling.
Figure 4 visualizes representative feature distributions and motivates combining attributes with relational structure.
Session node features:
  • Temporal and volume signals: session duration, request count, and request rate.
  • Coverage and depth: distinct pages/categories and indicators of multi-step actions (e.g., cart/login).
  • Lightweight fingerprints: coarse user-agent/headers when available; we avoid intrusive client-side telemetry.
Features are numeric/categorical (standardized/encoded) and kept lightweight for non-intrusive deployment.
Content (URL) node features: Each URL node has attributes describing page context, including:
  • Type/context: page category (e.g., product, search, home).
  • Global statistics: popularity/rarity and coarse sensitivity tags for special endpoints.
These URL attributes contextualize access patterns (e.g., concentrated browsing of rare or sensitive endpoints).
Motivation: Session-only models miss “feature-normal” bots; message passing combines behavior, page context, and shared neighborhoods.
III-C Model Architecture
We use GraphSAGE [2] to learn node representations via sampled neighbor aggregation. Its inductive formulation is critical because new sessions and URLs appear continuously.
GraphSAGE layers: Two layers capture 1-hop and 2-hop context. Each layer updates node $v$ as:
$\mathbf{h}_{v}^{(k)}=\sigma\left(W^{(k)}\cdot\text{AGG}\left(\mathbf{h}_{v}^{(k-1)},\mathbf{h}_{u_{1}}^{(k-1)},\mathbf{h}_{u_{2}}^{(k-1)},\ldots\right)\right),$ (1)
where $\mathbf{h}_{v}^{(0)}$ is the input feature vector, neighbors are $\{u_{1},u_{2},\ldots\}$, $\text{AGG}(\cdot)$ is permutation-invariant, $W^{(k)}$ are trainable weights, and $\sigma$ is a non-linearity (ReLU). We use a mean aggregator.
After two layers we obtain embeddings (e.g., 128-D) that summarize multi-hop neighborhoods: a session representation reflects its own features and the pages it visited, as well as other sessions that visited those pages.
Classifier: We apply an MLP head to session embeddings to output $P(\text{bot}\mid\text{session})$. Only session nodes are supervised; URL nodes participate in message passing as contextual carriers.
Inductiveness: GraphSAGE learns an aggregation function (not per-node embeddings), enabling scoring of new sessions/URLs from features and neighborhoods.
Why GraphSAGE: Neighbor sampling improves scalability on sparse, high-degree graphs, and the inductive formulation supports evolving graphs better than transductive alternatives.
III-D Training and Inference
Training: We train supervised on labeled sessions. Labels are obtained via a hybrid (semi-synthetic) strategy combining verified real-world attacks (e.g., honeypots/trap URLs) with controlled injections of diverse bot scripts; imbalance is handled via weighting/resampling.
We use binary cross-entropy on labeled session nodes:
$L=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right],$ (2)
where $y_{i}$ is the true label and $\hat{y}_{i}$ the predicted probability. We train with mini-batches by sampling target sessions and their $k$-hop neighborhoods (GraphSAGE sampling), with dropout and $L_{2}$ weight decay for regularization.
We select hyperparameters on validation AUC and recall under low false positive rates, reflecting operational constraints.
Inference: For a new session, we add its node/edges, compute features, aggregate a bounded neighborhood, and output a bot probability.
Inference runs on a bounded neighborhood subgraph, enabling near real-time scoring.
Incremental learning: The model can be periodically retrained with new confirmed samples and hot-swapped without changing online graph construction.
IV Experiments
We evaluate on real traffic logs and two simulations to test accuracy gains over a session-feature baseline and robustness to perturbations and unseen nodes.
IV-A Dataset and Experimental Setup
Dataset: We use anonymized server logs from a representative mid-sized e-commerce platform over two weeks, built via a hybrid (semi-synthetic) strategy for high-confidence labels. Background traffic consists of real production sessions ($\sim$80K) after removing sessions with <2<2 requests and truncating extreme outliers. Bots constitute a small fraction ($\sim$5%) and are drawn from verified real-world attacks (honeypots/trap URLs) as well as controlled injections (scrapers/headless browsers). The session–URL graph has on the order of $10^{5}$ edges (tens of thousands of sessions; thousands of URLs) with a power-law degree distribution; identifiers are anonymized/hashed. We use 10% validation and 10% test splits with similar class proportions, and a chronologically later test split to emulate deployment and evaluate inductive generalization.
Baselines: We compare against a session-feature MLP trained on the same session features but without graph connectivity. This isolates the value added by relational modeling.
We report AUC plus precision/recall/F1 at an operating threshold chosen to yield approximately 1% false positive rate on validation.
Training details: GraphSAGE uses two layers with 128-dimensional hidden states and neighbor sampling size 15; the MLP has two 128-unit hidden layers. Both use Adam (lr=0.001), early stopping on validation AUC, and class weights for imbalance. We run 5 seeds and report mean performance.
IV-B Performance Comparison
Table 1 shows that GraphSAGE (refined graph) outperforms the session-only MLP (AUC 0.9705 vs. 0.9102) and improves recall at $\sim$1% FPR. The raw-graph variant underperforms and is less stable, motivating refinement.
Table 1: Overall Bot Detection Performance Comparison on Test Set.
Model AUC (Mean $\pm$ Std) Precision Recall (@1% FPR) F1-Score
Session-level MLP (Baseline) 0.9102 $\pm$ 0.0150 0.7505 0.7510 0.7508
GraphSAGE (Raw Graph) 0.8756 $\pm$ 0.1042 0.8230 0.8105 0.8167
GraphSAGE (Ours, Refined) 0.9705 $\pm$ 0.0085 0.8055 0.9002 0.8501
Note: indicates results that align with the qualitative metrics reported in the deployment analysis.
11223344550.80.80.90.911Cross-Validation Fold IDAUC ScoreGraphSAGE (Refined)GraphSAGE (Raw)MLP (Session-only)
Figure 5: Fold-level AUC comparison across five cross-validation splits. Graph refinement substantially reduces performance variance compared to the raw graph, leading to more stable and reliable detection performance.
GraphSAGE recovers “feature-normal” bots that look benign in aggregates but exhibit atypical session–URL connectivity (e.g., rare-page mixtures), a signal absent from the MLP. Figure 5 reports fold-level stability.
IV-C Adversarial Perturbation Experiment
Bots may adapt their browsing to evade detection. We simulate adversarial perturbations by modifying session–URL edges; this complements injected bots by modeling lightweight evasive adaptations.
Setup: From the test graph, we perturb bot-session connectivity:
  • Edge addition: add a few edges from bot sessions to popular URLs to mimic “masking” via common page visits.
  • Edge removal: remove a few edges (typically to least popular pages) to mimic avoiding “red-flag” targets.
We vary intensity by edges modified per bot session and maintain feature consistency (e.g., request counts) under removals.
011223344550.70.70.80.80.90.911Perturbation Intensity (Edges Modified per Bot Session)AUC ScoreRobustness under Adversarial PerturbationGraphSAGE (Ours)MLP (Baseline)
Figure 6: Impact of adversarial graph perturbation on model performance. The session-level MLP degrades rapidly as adversarial edges disrupt feature consistency, while GraphSAGE exhibits substantially higher robustness under moderate perturbations (1–3 edges), benefiting from structural aggregation across session–URL interactions.
Results: Figure 6 shows modest degradation under mild perturbation (e.g., AUC 0.971 $\rightarrow$ 0.958 at 2 modified edges/session) while remaining above the MLP; heavier perturbations reduce the gap as bots mimic benign connectivity. Robustness benefits from combining structure and attributes, consistent with graph anomaly detection and robust graph learning [5, 7].
IV-D Cold-Start Simulation Experiment
Cold start is essential: new sessions and pages must be scored without retraining. We validate inductive generalization with a rolling simulation.
Setup: We train on week 1 and perform direct inductive inference on week 2, which introduces entirely new session nodes and some new URL nodes. We also report an optional fine-tuning upper bound and compare with the MLP baseline. No future labels or interactions are used during inductive inference.
Table 2: Generalization Capability: Cold-Start Simulation (Week 1 vs. Week 2).
Scenario Model Setting MLP (Baseline) GraphSAGE (Ours)
Week 1 (In-Sample) Trained on W1 0.9100 0.9705
Performance Gap - -
Week 2 (Cold-Start) Inductive Inference 0.8500 0.9630
Relative Drop $\downarrow 6.6\%$ $\downarrow 0.8\%$
Fine-tuned (Optional) N/A 0.9720
Direct inference on new session/URL nodes without retraining, demonstrating inductive generalization.
Results: GraphSAGE preserves strong performance under cold start (AUC 0.963 vs. 0.970 in-sample) with only a minor drop. The main failure mode is sparse context when sessions hit mostly unseen URLs. The MLP drops more under week-2 shift, while GraphSAGE remains stable; optional fine-tuning restores peak performance.
V Results and Discussion
We analyze where gains originate, baseline limitations, interpretability, and deployment considerations.
V-A Impact of Graph Structure vs. Semantic Features
We ablate topology vs. attributes by training (a) a structure-only GNN with minimal semantic features and (b) a feature-only model (MLP). The full model performs best, while the structure-only GNN still outperforms the feature-only baseline (AUC $\sim$0.88 vs. 0.85), confirming that topology carries critical signal and that topology and semantics are complementary in graph anomaly detection [5].
V-B Baseline MLP Performance and Limitations
The MLP baseline is strong, indicating session features capture obvious automation, but it fails most on feature-normal bots and under temporal shift. GraphSAGE mitigates these cases by incorporating relational context, consistent with graph-based fraud workflows [3].
V-C Model Interpretability and Case Study
Graph predictions can be inspected via a session’s local neighborhood (unique/rare URLs, shared target sets, or coordinated clusters), providing actionable explanations for analysts that are less transparent in feature-only models. For instance, we observed a bot cluster flagged primarily because its sessions shared access to an outdated API endpoint, a structural anomaly that is difficult to surface from session aggregates alone.
V-D Generalization and Adaptability
The perturbation and cold-start results indicate the model captures relationship-level signals that are harder to evade with small behavioral tweaks. More sophisticated coordination where each session appears benign remains a challenge and motivates richer graphs and higher-order pattern modeling.
V-E Deployment Considerations
The system deploys as a backend plug-in: feature extraction and graph updates feed a scoring service that outputs a bot risk score per session, using only existing logs. The model can be retrained offline and hot-swapped; inductive inference degrades gracefully on new patterns until retraining. We cap neighbor sampling to bound runtime; scoring a session with up to $\sim$50 page visits takes under 50 ms on CPU.
VI Conclusion
We presented a non-intrusive, graph-based bot detection framework evaluated on real-world traffic with high-confidence bot labels that models session–URL interactions and applies inductive GraphSAGE for session classification. By combining topology with semantic features, the refined-graph model improves over a strong session-only MLP while avoiding CAPTCHAs and client-side instrumentation. Robustness experiments show modest degradation under edge perturbations, and cold-start evaluation confirms strong inductive generalization to new sessions/URLs, supporting deployment as a bounded-latency backend scoring module.
VI-A Future Work
Future directions include richer heterogeneous graphs (e.g., adding account/IP nodes), explicit defenses against adversarial edge manipulation, and more fine-grained attribution to support analyst workflows, along with live A/B evaluation to quantify operational trade-offs.